# CtRL-Sim: Reactive and Controllable Driving Agents with Offline Reinforcement Learning

### v1

@@ 5 Related Work
Agent behaviour simulation involves modelling the behaviour of other agents in simulation, such as vehicles and pedestrians, to enable diverse and realistic interactions with the AV. Agent behaviour simulation methods can be categorized into rule-based and data driven methods. Rule based methods rely on human-specified rules to produce plausible agent behaviours, such as adhering strictly to the center of the lane and maintaining a safe following distance to a lead vehicle (Treiber et al., 2000; Kesting et al., 2007). These methods, while simple and interpretable, often yield unrealistic agent behaviours because they rigidly adhere to rules and thus fail to capture the full spectrum of human driving behaviours an AV might encounter in the real world. Moreover, we are most interested in modeling long-tail behaviours, which are difficult to model with rules alone.
@@
To address these limitations, prior work has proposed learning generative models that aim to replicate agent behaviours found in real-world driving trajectory datasets (Bergamini et al., 2021; Kamenev et al., 2022; Ścibior et al., 2021; Suo et al., 2021; Xu et al., 2023b; Wang et al., 2023; Philion et al., 2023). These approaches draw significant inspiration from the extensive array of methods proposed for the task of joint motion prediction (Casas et al., 2020; Girgis et al., 2021; Ngiam et al., 2022; Chen et al., 2022; Rowe et al., 2023); however, it’s crucial to distinguish that, unlike the open-loop nature of joint motion prediction, behavior simulation operates in a closed-loop manner (Caesar et al., 2021). To improve the realism of the learned behaviours, other work has proposed using adversarial imitation learning (Ho & Ermon, 2016) to minimize the behavioural discrepancy between expert and model rollouts (Bhattacharyya et al., 2018; Zheng et al., 2021; Igl et al., 2022) or RL to improve traffic rule compliance (Lu et al., 2023; Zhang et al., 2023a). While such methods demonstrate improved realism over rule-based methods, they lack the necessary control over the behaviours to enable the generation of targeted simulation scenarios for AV testing.
@@
More recent work has proposed more controllable behaviour simulation models by learning conditional models (Rempe et al., 2022; Zhang et al., 2023b; Suo et al., 2023; Chang et al., 2023; Zhong et al., 2023b; a; Chang et al., 2024) that enable conditioning on a high-level latent variables (Rempe et al., 2022; Zhang et al., 2023b), route information (Suo et al., 2023), or differentiable constraints (Zhong et al., 2023b; a; Xu et al., 2023a; Chang et al., 2024; Guo et al., 2023). More recently, Ding et al. (2023) used retrieval augmented generation to generate controllable traffic scenarios. However, these methods either lack interpretable control over the generated behaviours (Zhang et al., 2023b) or require costly test-time optimization procedures to steer the generated behaviours, such as latent variable optimization (Rempe et al., 2022), Bayesian optimization (Abeysirigoonawardena et al., 2019; Wang et al., 2021; Suo et al., 2023), or the simulation of expensive diffusion processes (Zhong et al., 2023b; Jiang et al., 2023; Zhong et al., 2023a; Xu et al., 2023a; Chang et al., 2024; Guo et al., 2023). CtRL-Sim takes an alternative approach and learns a conditional multi-agent behaviour model that conditions on interpretable factorized returns. By exponentially tilting the predicted return distribution (Lee et al., 2022) at test time, CtRL-Sim enables efficient, interpretable, and fine-grained control over agent behaviours while being grounded in real-world data.
@@ 6 Conclusions
Limitations. The learned policies produced by CtRL-Sim in its current form can be prohibitively slow for online reinforcement learning research as it requires a forward pass through a Transformer decoder at every simulation timestep. We leave the exploration of employing more lightweight policies through distillation methods for future work.
Acknowledgements
@@ Appendix A Offline RL Approaches
Table 3: Offline policy modelling approaches in prior work. We can see that methods differ in the decomposition of the joint distribution over actions and returns, with some approaches utilizing state prediction as a regularizer. We note that this table is adopted from prior work (Piché et al., 2022; Gontier et al., 2023).
@@ Appendix B Action Sampling Algorithm
Appendix C Nocturne Simulator
A scene in Nocturne is represented by a set of dynamic objects – such as vehicles, pedestrians, and cyclists – and the map context, which includes lane boundaries, lane markings, traffic signs, and crosswalks. Each dynamic object is prescribed a goal state, which is defined as the final waypoint in the ground-truth trajectory from the Waymo Open Motion Dataset. If there exist missing timesteps in the ground-truth trajectory, we re-define the goal as the waypoint immediately preceding the first missing timestep. By default, the dynamic objects track its 9 second trajectory from the Waymo Open Motion Dataset at 10 Hz. The Nocturne Simulator is originally designed for the development of RL driving policies, where the first 10 simulation steps (1s) of context is provided and the RL agent must reach the prescribed goal within the next 80 simulation steps (8s).
Appendix D Offline Reinforcement Learning Dataset
@@
$R_{g}(s_{t},s_{G})=\mathbbm{1}_{\text{goal achieved}}(s_{t},s_{G}).$ (6)
The vehicle-vehicle collision reward is defined by:
$R_{v}(s_{t},\mathbb{S}_{t}-\{s_{t}\})=-10\times\mathbbm{1}_{\text{vehicle-vehicle collision}}(s_{t},\mathbb{S}_{t}-\{s_{t}\})+\frac{\min(\text{dist-nearest-vehicle}(s_{t},\mathbb{S}_{t}-\{s_{t}\}),15)}{15},$ (7)
@@
Appendix E Evaluation Metrics
@@
where $m$ is the pointwise mean of $p$ and $q$ and $D_{\text{KL}}$ is the KL-divergence. Unlike prior works that compute the Jensen Shannon Divergence (Igl et al., 2022; Suo et al., 2023), we compute its square root – the Jensen Shannon Distance – so that values are not too close to 0. We compute the JSD over the following feature distributions: linear speed, angular speed, acceleration, and nearest distance. Since the acceleration values are discrete, for the acceleration JSD, we define one histogram bin for each valid acceleration value, yielding 21 evenly spaced bins between -10 and 10. For the linear speed histogram, we use 200 uniformly spaced bins between 0 and 30. For the angular speed JSD, we use 200 uniformly spaced bins between -50 and 50. For the nearest distance JSD, we use 200 uniformly spaced bins between 0 and 40.
Appendix F Individual JSD Results
@@
Actions-Only (Philion et al., 2023) 2.6 $\pm$ 0.3 18.1 $\pm$ 0.5 31.4 $\pm$ 0.5 5.3 $\pm$ 0.3 14.3 $\pm$ 0.4
Imitation Learning 1.1 $\pm$ 0.2 13.9 $\pm$ 0.3 31.9 $\pm$ 1.0 2.1 $\pm$ 0.2 12.2 $\pm$ 0.2
DT (Max Return) (Chen et al., 2021) 2.7 $\pm$ 0.2 13.6 $\pm$ 0.1 30.4 $\pm$ 0.6 2.1 $\pm$ 0.2 12.2 $\pm$ 0.2
CtRL-Sim (No State Prediction) 1.4 $\pm$ 0.1 14.2 $\pm$ 0.2 31.3 $\pm$ 0.8 2.1 $\pm$ 0.4 12.3 $\pm$ 0.3
CtRL-Sim (Base) 1.4 $\pm$ 0.1 14.4 $\pm$ 0.2 31.3 $\pm$ 0.3 2.1 $\pm$ 0.1 12.3 $\pm$ 0.1
CtRL-Sim (Positive Tilting) 1.6 $\pm$ 0.2 14.2 $\pm$ 0.1 30.9 $\pm$ 0.2 2.1 $\pm$ 0.1 12.2 $\pm$ 0.0
$\text{DT}^{*}$ (GT Initial Return) 1.3 $\pm$ 0.1 13.8 $\pm$ 0.2 31.5 $\pm$ 0.4 2.2 $\pm$ 0.1 12.2 $\pm$ 0.1
@@
Table 4: Breakdown of Meta-JSD in Table 1. For each metric, the best method is bolded and second-best method is underlined.
Appendix G CtRL-Sim Training Details
@@
The CtRL-Sim behaviour simulation model is trained using randomly subsampled sequences of length of $H\times N\times 3$, where $H=32$ and $N=24$. All agents and the map context are encoded in global frame as in (Ngiam et al., 2022; Philion et al., 2023) where we center and rotate the scene on a random agent during training. We select the $L=200$ closest lane segments within 100 metres of the centered agent as the map context, and select up to $N=24$ closest agents within 60 metres of the centered agent as social context for the model. For each lane segment, we subsample $P=100$ points. We use a hidden dimension size $d=256$, where we use $E=2$ Transformer encoder blocks and $D=4$ Transformer decoder blocks, and we set $\alpha=\frac{1}{100}$ in the loss function. We supervise our model only on the trajectories of moving agents. We found it useful to employ goal dropout whereby the embeddings for $10\%$ of agent goals are randomly set to 0 to prevent the model from overrelying on the goal information. We found goal dropout useful for learning an informative map representation. The state, return, and action embeddings for the missing timesteps are set to 0. The CtRL-Sim model is trained using a linear decaying learning rate schedule from 5e-4 for 200k steps using the AdamW optimizer and a batch size of 64. At inference, we sample actions with a temperature of 1.5. The CtRL-Sim architecture comprises 8.3 million parameters that we train in 20 hours with 4 NVIDIA A100 GPUs.
Appendix H CtRL-Sim Inference Details
@@
CtRL-Sim supports scenes with an arbitrary number of agents. As CtRL-Sim is trained with up to $N=24$ agents, when the number of CtRL-Sim-controlled agents at inference time exceeds $N=24$, we iteratively select 24-agent subsets at each timestep for processing until all agents have been processed. We first randomly select a CtRL-Sim-controlled agent, we normalize the scene to this agent and select the 23 closest context agents to the CtRL-Sim-controlled agent to comprise the first set of 24 agents. We then iteratively continue centering on a CtRL-Sim-controlled agent that has not been processed in the previous sets of 24 agents and select its 23 closest agents for context until all CtRL-Sim-controlled agents have been included in a 24-agent subset. If an agent belongs to multiple 24-agent subsets, we use the model’s first prediction of that agent. At inference time, the context length is set to training context length $H=32$. At each timestep, we select $H=32$ most recent timesteps as context and we found it useful to always center and rotate the scene on the centered agent at the oldest timestep in the context. For the first 10 timesteps (1s) of the simulated rollout, the states and actions are fixed to the ground-truth states and actions from the offline RL dataset, whereas the return-to-go is predicted at every timestep of the simulated rollout.
Appendix I CtRL-Sim Finetuning

### v2

@@ 4 Related Work
Agent behaviour simulation involves modelling the behaviour of other agents in simulation, such as vehicles and pedestrians, to enable diverse and realistic interactions with the AV. Agent behaviour simulation methods can be categorized into rule-based and data-driven methods. Rule-based methods rely on human-specified rules to produce plausible agent behaviours, such as adhering strictly to the center of the lane [2, 3]. These methods often yield unrealistic and rigid agent behaviours that fail to capture the full spectrum of driving behaviours. Moreover, we are most interested in modelling long-tail behaviours, which are difficult to model with rules alone.
@@
To address these limitations, prior work has proposed learning generative models that aim to replicate agent behaviours found in real-world driving trajectory datasets [22, 23, 24, 4, 5, 25, 9]. These approaches draw significant inspiration from the extensive array of methods proposed for the task of joint motion prediction [26, 27, 28, 29, 30]; however, it’s crucial to distinguish that, unlike the open-loop nature of joint motion prediction, behaviour simulation operates in a closed-loop manner [31]. To improve the realism of the learned behaviours, other work has proposed using adversarial imitation learning [32] to minimize the behavioural discrepancy between expert and model rollouts [33, 34, 6] or RL to improve traffic rule compliance [35, 36]. While such methods demonstrate improved realism over rule-based methods, they lack the necessary control over the behaviours to enable the generation of targeted simulation scenarios for AV testing.
@@
More recent work has proposed more controllable behaviour simulation models by learning conditional models [10, 37, 7, 38, 8, 11, 12] that enable conditioning on a high-level latent variables [10, 37], route information [7], or differentiable constraints [8, 11, 39, 12, 40]. More recently, [41] used retrieval augmented generation to generate controllable traffic scenarios. However, these methods either lack interpretable control over the generated behaviours [37] or require costly test-time optimization procedures to steer the generated behaviours, such as latent variable optimization [10], Bayesian optimization [42, 43, 7], or the simulation of expensive diffusion processes [8, 44, 11, 39, 12, 40]. CtRL-Sim takes an alternative approach and learns a conditional multi-agent behaviour model that conditions on interpretable factorized returns. By exponentially tilting the predicted return distribution [16] at test time, CtRL-Sim enables efficient, interpretable, and fine-grained control over agent behaviours while being grounded in real-world data.
@@ 5 Conclusions
Acknowledgments
@@ Appendix A Offline RL Approaches
Table 3: Offline policy modelling approaches in prior work. We can see that methods differ in the decomposition of the joint distribution over actions and returns, with some approaches utilizing state prediction as a regularizer. We note that this table is adopted from prior work [18, 17].
@@ Appendix B Action Sampling Algorithm
Appendix C Nocturne Physics Simulator and Offline RL Dataset
C.1 Physics-based Nocturne Simulator
CtRL-Sim extends the Nocturne simulation environment [13]. Nocturne is a lightweight 2D driving simulator that is built on real-world driving trajectory data from the Waymo Open Motion Dataset [14]. A scene in Nocturne is represented by a set of dynamic objects – such as vehicles, pedestrians, and cyclists – and the map context, which includes lane boundaries, lane markings, traffic signs, and crosswalks. Each dynamic object is prescribed a goal state, which is defined as the final waypoint in the ground-truth trajectory from the Waymo Open Motion Dataset. If there exist missing timesteps in the ground-truth trajectory, we re-define the goal as the waypoint immediately preceding the first missing timestep. By default, the dynamic objects track its 9 second trajectory from the Waymo Open Motion Dataset at 10 Hz. The Nocturne Simulator is originally designed for the development of RL driving policies, where the first 10 simulation steps (1s) of context is provided and the RL agent must reach the prescribed goal within the next 80 simulation steps (8s).
We extend Nocturne by integrating a physics engine based on the Box2D library for enabling realistic vehicle dynamics and vehicle collisions. We model the vehicle’s dynamics using basic physics principles, where forces applied to the vehicle are translated into acceleration, influencing its speed and direction, and with frictional forces applied to simulate realistic sliding and adherence behaviors. This extension additionally ensures that an agent’s acceleration, braking, and turn radius are bound by plausible limits and that vehicles can physically collide with each other. Such improvements open the possibility of more accurately simulating complex conditions, such as emergency braking maneuvers, slippery roads, and multi-vehicle collisions.
C.2 Offline RL Dataset Collection
The actions are defined by the acceleration and steering angle and the reward function is decomposed into three components: a goal position reward, a vehicle to vehicle collision reward, and a vehicle to road edge collision reward. We confirm that the trajectory rollouts obtained by feeding Waymo scenes through the simulator attain a reasonable reconstruction of the ground-truth Waymo trajectories (see Table 1). Following Nocturne [13], we omit bicyclist and pedestrian trajectories from the Waymo Open Motion Dataset and we omit scenes containing traffic lights. This yields a training, validation, and test set containing 134150, 9678, and 2492 scenes.
@@ C.2 Offline RL Dataset Collection
$R_{g}(s_{t},s_{G})=\mathbbm{1}_{\text{goal achieved}}(s_{t},s_{G}),$ (2)
[…]
$R_{v}(s_{t},\mathbb{S}_{t}-\{s_{t}\})$ $=-10\times\mathbbm{1}_{\text{vehicle-vehicle collision}}(s_{t},\mathbb{S}_{t}-\{s_{t}\})$ (3)
$+\frac{\min(\text{dist-nearest-vehicle}(s_{t},\mathbb{S}_{t}-\{s_{t}\}),15)}{15},$ (4)
@@
Appendix D Evaluation Metrics
@@
where $m$ is the pointwise mean of $p$ and $q$ and $D_{\text{KL}}$ is the KL-divergence. Unlike prior works that compute the Jensen Shannon Divergence [6, 7], we compute its square root – the Jensen Shannon Distance – so that values are not too close to 0. We compute the JSD over the following feature distributions: linear speed, angular speed, acceleration, and nearest distance. Since the acceleration values are discrete, for the acceleration JSD, we define one histogram bin for each valid acceleration value, yielding 21 evenly spaced bins between -10 and 10. For the linear speed histogram, we use 200 uniformly spaced bins between 0 and 30. For the angular speed JSD, we use 200 uniformly spaced bins between -50 and 50. For the nearest distance JSD, we use 200 uniformly spaced bins between 0 and 40.
Appendix E Individual JSD Results
@@
Actions-Only [9] 4.1 $\pm$ 0.7 16.8 $\pm$ 0.3 15.6 $\pm$ 0.5 5.1 $\pm$ 0.5 10.4 $\pm$ 0.3
Imitation Learning 1.0 $\pm$ 0.1 13.4 $\pm$ 0.3 16.9 $\pm$ 0.4 2.1 $\pm$ 0.2 8.3 $\pm$ 0.1
DT (Max Return) [15] 2.7 $\pm$ 0.1 13.4 $\pm$ 0.2 15.4 $\pm$ 0.5 2.2 $\pm$ 0.3 8.4 $\pm$ 0.1
CTG++ [11] 3.2 $\pm$ 0.9 11.9 $\pm$ 0.9 12.4 $\pm$ 0.4 2.2 $\pm$ 0.3 7.4 $\pm$ 0.2
CtRL-Sim (No State Prediction) 1.2 $\pm$ 0.1 13.7 $\pm$ 0.2 15.9 $\pm$ 0.7 2.0 $\pm$ 0.2 8.2 $\pm$ 0.2
CtRL-Sim (Base) 1.1 $\pm$ 0.2 13.8 $\pm$ 0.2 15.6 $\pm$ 0.5 2.0 $\pm$ 0.3 8.1 $\pm$ 0.2
CtRL-Sim (Positive Tilting) 1.4 $\pm$ 0.1 13.6 $\pm$ 0.2 14.8 $\pm$ 0.5 1.8 $\pm$ 0.2 7.9 $\pm$ 0.1
$\text{DT}^{*}$ (GT Initial Return) 1.1 $\pm$ 0.2 13.4 $\pm$ 0.2 16.8 $\pm$ 0.6 2.1 $\pm$ 0.2 8.4 $\pm$ 0.1
@@
Table 4: Breakdown of Meta-JSD in Table 1. For each metric, the best unprivileged method is bolded and second-best is underlined. denotes a privileged method requiring the ground-truth future trajectory.
Appendix F CtRL-Sim Training and Inference Details
@@
Training: The CtRL-Sim behaviour simulation model is trained using randomly subsampled sequences of length of $H\times N\times 3$, where $H=32$ and $N=24$. For the actions, we discretize the acceleration and steering into 20 and 50 uniformly quantized bins, respectively, yielding 1000 action tokens. For the return-to-gos, we discretize each return-to-go component $G^{c,i}_{t}$ into 350 uniformly quantized bins. All agents and the map context are encoded in global frame as in [28, 9] where we center and rotate the scene on a random agent during training. The map context is represented as a set of road segments $m:=\{\mathbf{r}_{l}\}_{l=1}^{L}$, where each road segment is defined by a sequence of points $\mathbf{r}_{l}:=(p_{l}^{1},\dots p_{l}^{P})$, where $L$ is the number of road segments and $P$ is the number of points per road segment. We apply a per-point MLP to the points of each road segment $\mathbf{r}_{l}$. To produce road segment-level embeddings, we then apply attention-based pooling [45] on the embeddings of the points within each road segment, yielding $L$ road segment embeddings of size $d$. We select the $L=200$ closest lane segments within 100 metres of the centered agent as the map context, and select up to $N=24$ closest agents within 60 metres of the centered agent as social context for the model. For each lane segment, we subsample $P=100$ points. We use a hidden dimension size $d=256$, where we use $E=2$ Transformer encoder blocks and $D=4$ Transformer decoder blocks, and we set $\alpha=\frac{1}{100}$ in the loss function. We supervise our model only on the trajectories of moving agents. We found it useful to employ goal dropout whereby the embeddings for $10\%$ of agent goals are randomly set to 0 to prevent the model from overrelying on the goal information. We found goal dropout useful for learning an informative map representation. The state, return, and action embeddings for the missing timesteps are set to 0. To ensure that the model is permutation equivariant to the agent ordering [28, 27], we modify the standard temporally causal mask by additionally enforcing that each agent can only attend to its own action and return-to-go tokens at the present timestep while allowing access to all agents’ state tokens at the present timestep and all agents’ tokens in the past timesteps. The CtRL-Sim model is trained using a linear decaying learning rate schedule from 5e-4 for 200k steps using the AdamW optimizer and a batch size of 64. At inference, we sample actions with a temperature of 1.5. The CtRL-Sim architecture comprises 8.3 million parameters that we train in 20 hours with 4 NVIDIA A100 GPUs.
@@
Inference: CtRL-Sim supports scenes with an arbitrary number of agents. As CtRL-Sim is trained with up to $N=24$ agents, when the number of CtRL-Sim-controlled agents at inference time exceeds $N=24$, we iteratively select 24-agent subsets at each timestep for processing until all agents have been processed. We first randomly select a CtRL-Sim-controlled agent, we normalize the scene to this agent and select the 23 closest context agents to the CtRL-Sim-controlled agent to comprise the first set of 24 agents. We then iteratively continue centering on a CtRL-Sim-controlled agent that has not been processed in the previous sets of 24 agents and select its 23 closest agents for context until all CtRL-Sim-controlled agents have been included in a 24-agent subset. If an agent belongs to multiple 24-agent subsets, we use the model’s first prediction of that agent. At inference time, the context length is set to training context length $H=32$. At each timestep, we select $H=32$ most recent timesteps as context and we found it useful to always center and rotate the scene on the centered agent at the oldest timestep in the context. For the first 10 timesteps (1s) of the simulated rollout, the states and actions are fixed to the ground-truth states and actions from the offline RL dataset, whereas the return-to-go is predicted at every timestep of the simulated rollout.
Appendix G Baseline Details
In this section, we describe the design decision of each baseline employed in our work. We note that for all models below, we scaled them in order for all models to have approximately the same number of learnable parameters as CtRL-Sim’s architecture.
Actions Only
The actions-only baseline is encoder-decoder architecture implemented in exactly the same way as CtRL-Sim with a few ablations. These include removal of states and returns from the decoder sequence, and no state rollout predictions. This model was inspired by [9] but differs in that the model also has access to the agents’ goals.
Imitation Learning
The imitation learning baseline is also based on the CtRL-Sim multi-agent behaviour simulation architecture but lacks factorized return information. It is a step better than the actions-only baseline since it considers the states in the decoder sequence, and this is corroborated by its improved performance on multi-agent simulation results of Table 1.
DT
The Decision Transformer baseline is based on the seminal work [15]. We adopt an identical architecture to CtRL-Sim’s with some minor difference based on the algorithm. One such decision is the lack of a return prediction based on states, and instead returns are chosen at inference time based on domain knowledge. Returns are the first token fed to the decoder, followed by states in order to predict actions. We make the strong argument that this is suboptimal for controllability (results in Figure 4) and does not provide intuitive mechanism for selecting the return values to target.
CTG++
The CTG++ baseline is a diffusion model reimplementation of the recent work by [11]. We attempted to follow the architecture as closely as possible with a few minor differences. One such difference is that we diffuse over both states and actions, rather than diffusing over only actions and using an unicycle dynamics model to derive the states. We chose this approach because the underlying dynamics of the physics-enhanced Nocturne simulator is not necessarily governed by a unicycle dynamics model, and thus using a unicycle dynamics model would induce small errors in the derived states during training. We further note that we cannot replace the unicycle dynamics model with the Nocturne physics dynamics model as this forward model is not differentiable. We also condition on the present timestep and goal, to ensure fair comparison with CtRL-Sim. We note that at scene generation time, we diffuse over actions and states at a rate of 2 Hz which is consistent with the original CTG++ model. In addition, although we train with 100 diffusion steps, we run evaluations with 50 diffusion steps. As showing in Table 5, the difference in performance is insignificant. As we do not have information on the size of the network used in the original manuscript, we explored different hidden dimension sizes of the transformer architecture of the diffusion model, also shown in Table 5. A final difference is the future relative encoding. While CTG++ use the ground-truth to compute the relative encoding during training and a constant velocity model at test time, we opted to use the final historical timestep’s relative encoding. We found this approach to be more stable.
ADE FDE Goal Success JSD Collision Off Road Per Scene
Method (m) $\downarrow$ (m) $\downarrow$ Rate (%) $\uparrow$ (×$\times 10^{-2}$) $\downarrow$ (%) $\downarrow$ (%) $\downarrow$ Gen. Time (s) $\downarrow$
CTG++ $1.72$ $3.97$ $41.7$ $7.7$ $6.4$ $17.4$ $44.0$
CTG++ (128 hidden dim) $1.83$ $4.32$ $37.8$ $7.6$ $7.0$ $17.7$ $25.0$
CTG++ (100 diffusion steps) $1.87$ $3.98$ $43.0$ $8.6$ $7.9$ $16.8$ $140.0$
Table 5: Ablations of CTG++ on Multi-agent simulation results over 1000 test scenes. We report the results across all metrics of different configurations of the CTG++ baseline [11]. $\dagger$ indicates the original model results, with $256$ hidden dimension and $50$ diffusion steps at inference time.
Appendix H CAT Simulated Data Collection and CtRL-Sim Finetuning
The Waymo Open Motion dataset largely contains nominal driving scenes. To enhance control over the generation of safety-critical scenarios, we finetune CtRL-Sim on a simulated dataset of safety-critical scenarios generated by CAT [21]. CAT is a state-of-the-art collision generation method that involves fixing the agent’s future trajectory to a trajectory predicted by a DenseTNT trajectory predictor. CAT searches for a trajectory that has high likelihood of colliding with the log-replay future trajectory of the ego vehicle, while having high probability under the behaviour prior (DenseTNT). For more details, we refer readers to [21]. We note that a limitation of CAT is that the agent is non-reactive to the ego as the agent’s trajectory is fixed at the beginning of the simulation. Moreover, unlike CtRL-Sim, CAT does not have control over the degree to which the agent is adversarial.
To collect the simulated safety-critical dataset, we run CAT on a subset of the interactive validation split of the Waymo Open Motion Dataset, which involves two interacting agents. Following CAT, we select one of the two interacting agents to be the ego (whose trajectory is fixed to the log-replay trajectory) and the other interacting agent to be the CAT adversary. In total, we collect 3577 CAT scenarios for finetuning, of which around 60% contain ego-adversary collisions.
Method Times Preferred
CtRL-Sim 15
CAT 8
Tie 2
Table 6: The tally of votes for the pilot study. We show the breakdown of the votes for the conducted pilot study. Strong preference in terms of plausibility is shown for scenarios generated using CtRL-Sim.